# Automated Privacy Information Annotation in Large Language Model Interactions

### v1

@@ Abstract
Users interacting with large language models (LLMs) under their real identifiers often unknowingly risk disclosing private information. Automatically notifying users whether their queries leak privacy and which phrases leak what private information has therefore become a practical need. Existing privacy detection methods, however, were designed for different objectives and application scenarios, typically tagging personally identifiable information (PII) in anonymous content. In this work, to support the development and evaluation of privacy detection models for LLM interactions that are deployable on local user devices, we construct a large-scale multilingual dataset with 249K user queries and 154K annotated privacy phrases. In particular, we build an automated privacy annotation pipeline with cloud-based strong LLMs to automatically extract privacy phrases from dialogue datasets and annotate leaked information. We also design evaluation metrics at the levels of privacy leakage, extracted privacy phrase, and privacy information. We further establish baseline methods using light-weight LLMs with both tuning-free and tuning-based methods, and report a comprehensive evaluation of their performance. Evaluation results reveal a gap between current performance and the requirements of real-world LLM applications, motivating future research into more effective local privacy detection methods grounded in our dataset [^1] The code are public at https://github.com/Nidryen-zh/PrivacyAnnotation and the dataset can be found at https://huggingface.co/datasets/Nidhogg-zh/Interaction__\__Dialogue__\__with__\__Privacy..
1 Introduction
Figure 1: An example of privacy exposure.
@@ 1 Introduction
Large language models (LLMs) have demonstrated immense potential in revolutionizing user interactions with artificial intelligence. However, during cloud-based LLM interactions, users often unintentionally expose privacy to service providers privacyconcern ; survey ; mope ; assessingprivacy ; LPLM ; beyond ; decoding . A case study of the ShareGPT corpus sharegpt reveals that nearly one-third of dialogues contain privacy leaks. One such instance of exposing the user’s home address is illustrated in Figure 1. Even more concerning, LLM service providers typically require access to users’ or organizations’ identity information, exposing unique identifiers before any interaction occurs. This real-identity interaction setting significantly increases the risk of privacy information disclosure. Therefore, proactively alerting users to potential privacy exposure during interactions with LLMs has become an urgent and practical need. When a user launches a query, whether the query contains private information or not should be informed. If so, the specific phrases in the query responsible for the privacy leakage need to be annotated. To further assist users who may still be uncertain, a concise explanation of what private information will be exposed by each annotated phrase should be provided. These notifications can help users quickly understand the nature of privacy leakage and take appropriate action, such as rewriting the query.
@@
Previous research across different domains, including medical privacy medical1 ; medical2 ; medical3 , code privacy finding , and employment information Multilingual ; occupational , primarily focuses on detecting privacy risks of revealing individuals’ identities in anonymous online content or documents selfdisclosures ; entity2 ; synthetic . However, these scenarios differ significantly from privacy detection in LLM interactions. By authenticating identifiers, users pose queries to LLMs with their real names. It is important to detect privacy information related to the current users from these queries. If users disclose private information about unrelated third parties, such information will not be extracted since it does not constitute a harmful breach of privacy for the current user. Additionally, many existing studies are constrained to a limited set of topics or categories ner1 ; anonymisation ; categorylimit1 ; categorylimit2 . In contrast, privacy information in LLM interactions is much more diverse and cannot easily be classified into a limited set of predefined types.
@@
To enable privacy detection for real-time user queries, it is desired to deploy an efficient model on the user side for automatic local detection before sending queries to cloud-based LLMs. The development and evaluation of such a model require large-scale datasets to ensure good generalization ability. However, there does not exist privacy annotated datasets for user interaction dialogues. Existing datasets were manually labeled and limited to particular domains or categories, with only a few thousand entries tab ; selfdisclosures ; i2b2 ; cegs ; codealltag . To address this, we design an automated scalable privacy annotation pipeline to extract privacy phrases and summarize the corresponding privacy information for user dialogues. With the assistance of GPT-4, our pipeline operates through four systematic steps: (1) privacy leakage classification, which filters the input to determine whether the user’s query leaks privacy; (2) extensive privacy categories extraction, which identifies privacy categories to guide the extraction of phrases; (3) privacy phrase extraction, where privacy phrases are extracted from queries and filtered; and (4) privacy information annotation, which summarizes details for each phrase. Based on the automated pipeline, we extract privacy phrases and annotate privacy information for dialogue datasets in different languages, containing 97,659 user queries, 85,320 phrases and corresponding information for English portion and 151,988 user queries, 68,910 phrases and information for Chinese portion. We also conduct human evaluation, demonstrating that the results obtained through our pipeline are highly consistent with those manually annotated. Compared to directly prompting LLM to extract privacy phrases ghost , our pipeline produces more accurate and diverse annotations.
@@
The constructed dataset can serve not only as a benchmark for evaluation but also as a valuable source of training data for local privacy detection methods. For evaluation, we propose new metrics to privacy detection performance at the query, phrase, and information levels. We also set up tuning-free baselines, including prompting and in-context learning with locally deployable, light-weight LLMs, as well as supervised fine-tuning baselines. Benchmark results show that the fine-tuned 1B model with our dataset outperforms the directly prompted 72B model.
@@
We summarize the key contributions as follows: (1) We identify a new application requirement of privacy detection for real-name user interactions with LLMs; (2) we build the automated privacy annotation pipeline to classify privacy leakage, extract privacy phrase and information, resulting in the first large-scale dataset for privacy detection in real-name interaction dialogues. We also design new metrics for benchmark testing; and (3) we report the performance of local privacy detection baselines, which, despite their progress, still fall short of meeting practical requirements. This highlights the need for further research into more effective methods grounded in our dataset. Specifically, the best-performing fine-tuned model achieves 87.6% accuracy in privacy leakage classification, a 74.3% F1 score in privacy phrase extraction, and a 44.7% F1 score in privacy information summarization.
2 Related Work
@@
Privacy Risks of LLMs. Recent work constructs the corpus with online personal information or user interaction history privacysurvey2 ; finetuneingleak to train LLMs characterllm ; exploitingpersona , and enhance inference ability learningretrieval ; hallucination , resulting in the issue of memorization memorization ; privacysurvey1 ; ProPILE ; extractingtraining , which recalls the personal sequences from LLM corpus. Existing works solve this problem on the LLM service provider side, including data pre-processing approaches (e.g., sanitizing training data sanitizing and reducing the frequency of private data appearing Deduplicating1 ; Deduplicating2 ), and privacy-preserving training methods (e.g., differentially private training differential1 ; differential2 ; differential3 ). In contrast, we takes a different angle for privacy risks of LLMs by locally detecting privacy in the users’ interactions to prevent private information uploaded to LLM service providers, solving the problem at the root of the the user side.
@@
Privacy Detection. Most previous work focuses on privacy detection for specific documents or context shared by users on the Internet semantics ; privacypreserving , detecting privacy that can identify an individual for anonymization entity1 ; entity2 ; synthetic . However, privacy in user interaction dialogues is fundamentally different from these scenarios. During the interaction, users have already authenticated their identifiers. The target application scenario of this work is to detect any newly leaked private data related to the user after a query is made, extending beyond just identifiable information. Existing approaches are also typically constrained to specific topics or categories anonymisation ; personalinformation , such as medical privacy medical1 ; medical2 ; medical3 ; medical4 , personal employment information Multilingual ; occupational , online comments selfdisclosures , and code privacy finding . In contrast, the types of privacy information in user interaction is not limited. Furthermore, most existing privacy detection models are based on discriminative models, treating the detection process as a sequence tagging task selfdisclosures ; ner1 ; entity3 . They can not generate specific privacy information to remind user. MemoAnalyzer ghost directly prompt generative LLMs to identify open privacy information without predefining categories, but the detection performance remains suboptimal.
@@
Privacy Datasets. Existing datasets for privacy detection are limited to specific domains and there is no large-scale dataset for real-name user interaction dialogues. For instance, PrivacyBot privacybot provides a sentence-level privacy classification dataset, while I2B2/UTHealth i2b2 and CEGS N-GRID cegs focus on privacy entity recognition in the personal health domain. CodE ALLTAG codealltag targets privacy entity recognition in German emails, TAB tab annotates privacy entities in court judgments, and JobStack jobstack specializes in privacy entity annotation in job postings. Additionally, these datasets face challenges of data volume and they are manually labeled. In contrast, our dataset focuses on LLM interactions and is annotated automatically, resulting in significantly larger scale compared to existing datasets and encompassing a broader range of privacy. Furthermore, it offers detailed information for each extracted privacy phrase. The comparison details are provided in Appendix D.
3 Problem Formulation
@@
Privacy in LLM Interactions. We focus on the scenario where the user interacts with LLMs by submitting real-name text queries. The LLM service providers are honest-but-curious, which implies that they may attempt to infer privacy information from queries and do not actively launch attacks. Different from PII, which refers to data that can identify an individual in an anonymous text, we define privacy information in LLM interactions as the service provider’s knowledge about user-related private data or personal details, before and after the user poses a query with real unique identifier. The definition of privacy information leakage is consistent with the principle of differential privacy, which considers the difference of the query output with and without each individual data record. The privacy information in real-name user dialogues is regarded as the information, if exposed, will harm the personal, social, or professional aspects of the user who poses the query or the user’s related
@@
Figure 2: Automated privacy annotation pipeline with LLM API and privacy detection module with light-weight LLM.
people, such as family, friends, or close associates. Typical privacy information includes, but is not limited to, preferences, intentions, and opinions. We emphasize that the privacy information in LLM interactions should be related to the user who pose the query with direct evidences. For example, as shown in Figure 1, the statement “Shanghai is also where David works.” leaks PII about David, but it does not leaks the privacy information about the user, since there is no direct evidence indicating David is related to the user.
@@
Automated Privacy Annotation Requirement. As shown in Figure 2, it is necessary to first build an automated privacy annotation pipeline based on a cloud-side LLM API to annotate dialogue datasets with privacy phrases and specific privacy information. Such a dataset can serve as an evaluation benchmark for downstream applications and provide training data for optimization. Local privacy detection methods with light-weight LLM are then designed to detect privacy from user real-time queries before they are sent to service providers and they should be deployed locally on the user’s device or be a request proxy built by the user’s organization. Formally, we let $D$ denote the user dialogue and $q$ denote a query from $D$. The goal is to extract the privacy phrase $\mathcal{P}_{q}=\{p_{1},...,p_{m}\}$ and summarize specific privacy information $\mathcal{I}_{q}=\{I_{1},...,I_{m}\}$, where $m$ is the number of the extracted phrases, $p_{j}$ is defined as a segment of $q$ that discloses private or sensitive information about the user, and $I_{j}$ is a statement summarizing the specific details revealed by $p_{j}$.
4 Automated Privacy Annotation Pipeline for Dialogue Datasets
We design an automated four-step pipeline based on strong LLM API to extract privacy phrase and annotate privacy information for dialogue datasets. The comparison between the results of our automated pipeline and manual annotation demonstrates a high degree of consistency.
4.1 Pipeline Rationale
A straightforward approach to extract privacy information is directly prompting GPT-4. However, direct prompting may force GPT-4 to extract information from samples that do not contain any privacy leaks, which results in a decrease in precision. Furthermore, the original understanding of privacy by GPT-4 revolves around the idea of whether a privacy phrase can reveal a user’s identifiable information. While this is intuitive for identifying PII, it does not align with our intended focus in Section 3, which aims to detect privacy information in real-name user interactions. We also observe that the results of direct prompting are limited to a few common types of privacy, such as location, name, and gender, with less accuracy in detecting more nuanced aspects like user status or intent.
To address these limitations, we developed the privacy annotation pipeline, as shown in Figure 3. (1) To avoid privacy extraction for user dialogue that do not involve privacy leakage, the first step is a filter, classifying dialogues into two categories: leaked and non-leaked. (2) The second step aims for comprehensive coverage of privacy categories. We independently extract privacy categories for each sample and take the union of all categories as the extensive set of privacy categories at the dataset level. (3) Then, with the assistant of the extensive category set, we extract privacy phrase for each sample based on the whole context of queries and filter the results based on the rules of being user-related and having clear reference. (4) The last step is to annotate privacy information for each extracted phrase. We also compare the performance of direct prompting with our automated pipeline in Section 4.3. All associated prompts can be found in Appendix A.
4.2 Pipeline Design
Privacy Leakage or Not Classification. We explicitly address the scenario where user interacts with identifier exposure and prompt GPT-4 to assess whether the user has disclosed any new privacy information during the interaction. We further clarify that privacy information must be linked to the user or individuals associated with them, and it encompasses various aspects such as opinions, preferences, and personal details. To ensure more reliable inferences, we not only ask GPT-4 to generate a simple yes or no judgment, but also request it to explain the reasoning behind its decision. Additionally, we provide a set of manually annotated samples as reference to guide the model.
@@ 4.2 Pipeline Design
Figure 3: Overview of our automated pipeline to extract privacy phrase and annotate privacy information over dialogue datasets: ➊ determining whether the input leaks privacy; ➋ extracting privacy categories from user queries that leaks privacy; ➌ extracting privacy phrases from user queries with the help of privacy categories; and ➍ annotating privacy information for each phrase.

### v2

# Automated Annotation of Privacy Information in User Interactions with Large Language Models

@@ Abstract.
Users interacting with large language models (LLMs) under their real identifiers often unknowingly risk disclosing private information. Automatically notifying users whether their queries leak privacy and which phrases leak what private information has therefore become a practical need. Existing privacy detection methods, however, were designed for different objectives and application domains, typically tagging personally identifiable information (PII) in anonymous content, which is insufficient in real-name interaction scenarios with LLMs. In this work, to support the development and evaluation of privacy detection models for LLM interactions that are deployable on local user devices, we construct a large-scale multilingual dataset with 249K user queries and 154K annotated privacy phrases. In particular, we build an automated privacy annotation pipeline with strong LLMs to automatically extract privacy phrases from dialogue datasets and annotate leaked information. We also design evaluation metrics at the levels of privacy leakage, extracted privacy phrase, and privacy information. We further establish baseline methods using light-weight LLMs with both tuning-free and tuning-based methods, and report a comprehensive evaluation of their performance. Evaluation results reveal a gap between current performance and the requirements of real-world LLM applications, motivating future research into more effective local privacy detection methods grounded in our dataset [^1] The code and dataset are public at https://github.com/Nidryen-zh/PrivacyAnnotation and https://huggingface.co/datasets/Nidhogg-zh/Interaction_\_Dialogue_\_with_\_Privacy..
LLM Interaction, Privacy Detection Dataset
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: In Proceedings of the 32th ACM SIGKDD Conference on Knowledge Discovery and Data Mining ; August 9-13, 2026; Jeju, Koreaisbn: 978-1-4503-XXXX-X/2025/08ccs: Security and privacyccs: Security and privacy Human and societal aspects of security and privacyccs: Security and privacy Privacy protectionsccs: Computing methodologiesccs: Computing methodologies Natural language processingccs: Computing methodologies Discourse, dialogue and pragmatics
1. Introduction
@@
Large language models (LLMs) have demonstrated immense potential in revolutionizing user interactions with artificial intelligence. However, during cloud-based LLM interactions, users often unintentionally expose privacy to service providers (Saglam et al., 2021; Yao et al., 2023; Li et al., 2023; Tang et al., 2023; Huang et al., 2022; Staab et al., 2024; Falade, 2023; Mireshghallah et al., 2024). A case study of the ShareGPT corpus (ShareGPT, 2023) reveals that nearly one-third of dialogues contain privacy leaks. One such instance of exposing the user’s intents and preferences is illustrated in Figure 1. Even more concerning, LLM service providers typically require access to users’ or organizations’ identity information, exposing unique identifiers before any interaction occurs. This real-identity interaction setting significantly increases the risk of privacy information disclosure. Therefore, proactively alerting users to potential privacy exposure during interactions with LLMs has become an urgent and practical need. When a user launches a query, whether the query contains private information or not, should be informed. If so, the specific phrases in the query responsible for the privacy leakage need to be annotated. To further assist users who may still be uncertain, a concise explanation of what private information will be exposed by each annotated phrase should be provided. These notifications can help users quickly understand the nature of privacy leakage and take appropriate action, such as rewriting the query.
Figure 1. An illustration of privacy exposure and annotation during a user’s interactions with LLM.
@@
Previous research for privacy protection across different domains, including medical privacy (Zhao et al., 2019; Klein et al., 2017; Valizadeh et al., 2023), code privacy (Tang and Østvold, 2024), and employment information (Tonneau et al., 2022; Preotiuc-Pietro et al., 2015), primarily focuses on detecting privacy risks by identifying private entities that reveals personal identifiable information (PII) in anonymous online content or documents (Dou et al., 2024; Papadopoulou et al., 2022; Yukhymenko et al., 2024a). However, these scenarios differ significantly from privacy detection in LLM interactions. By authenticating through user login, queries posed to LLMs are linked to a specific user’s real identifier. It is important to detect privacy information related to the current user from queries. On the other hand, if users disclose private information about unrelated third parties, such information will not be extracted since it does not constitute a harmful breach of privacy for the current user. Additionally, many existing studies are constrained to a limited set of topics or categories (Subramani et al., 2023; Lison et al., 2021a; Lee et al., 2023; Reuel et al., 2022). In contrast, privacy information in LLM interactions is much more diverse and cannot easily be classified into a limited set of types.
@@
To enable privacy detection for real-time user queries during the interactions, it is desired to deploy an efficient model on the user side for automatic local detection before sending queries to cloud-based LLMs. The development and evaluation of such a model require large-scale datasets to ensure good generalization ability. However, no such comprehensive datasets currently exist for LLM interactions. Existing datasets were manually labeled and limited to particular domains or categories, with only a few thousand entries (Pilán et al., 2022; Dou et al., 2024; Stubbs and Uzuner, 2015; Filannino et al., 2018; Eder et al., 2020). To address this, we design an automated, scalable privacy annotation pipeline to extract privacy phrases and summarize the corresponding privacy information for user dialogues. With the assistance of GPT-4o, our pipeline operates through four systematic steps: (1) privacy leakage classification, which filters the input to determine whether the user’s query leaks privacy; (2) extensive privacy categories extraction, which identifies privacy categories to guide the extraction of phrases; (3) privacy phrase extraction, where privacy phrases are extracted from queries and filtered; and (4) privacy information annotation, which summarizes details for each privacy phrase. Based on the automated pipeline, we extract privacy phrases and annotate privacy information for dialogue datasets in different languages, containing 97,659 user queries, 85,320 phrases with corresponding information for the English portion, and 151,988 user queries, 68,910 phrases with information for the Chinese portion. We also conduct human evaluation, demonstrating that the results obtained through our pipeline are highly consistent with those manually annotated. Compared to directly prompting LLM to extract privacy phrases (Zhang et al., 2024), our pipeline produces more accurate and diverse annotations.
@@
The constructed dataset can serve not only as a benchmark for evaluation but also as a valuable source of training data for local privacy detection methods. For evaluation, we design a hierarchical set of metrics at query, phrase, and information levels, each corresponding to tasks of increasing granularity in privacy detection. Specifically, we use accuracy for privacy leakage classification, which determines whether a query reveals private information. For privacy phrase extraction, which identifies sensitive phrases within queries, and privacy information summarization, which captures the specific private information leaked, we design precision, recall, and F1 scores to assess performance. These metrics quantify the differences between the extracted phrases or generated privacy information and the ground truth, effectively evaluating both missed and incorrect extractions or summaries. We also set up tuning-free baselines, including prompting and in-context learning with locally deployable, light-weight LLMs, as well as supervised fine-tuning baselines. Benchmark results show that the fine-tuned 1B model with our dataset outperforms the directly prompted 72B model.
@@
We summarize the key contributions as follows: (1) We identify a new application requirement of privacy detection for real-name user interactions with LLMs; (2) we build the automated privacy annotation pipeline to classify privacy leakage, extract privacy phrases and information, resulting in the first large-scale dataset for privacy detection in real-name interactions. We also design new metrics for benchmark testing; and (3) we report the performance of local privacy detection baselines, which, despite their progress, still fall short of meeting practical requirements. This highlights the need for further research into more effective methods grounded in our dataset. Specifically, the best-performing fine-tuned model achieves 87.6% accuracy in privacy leakage classification, a 74.3% F1 score in privacy phrase extraction, and a 44.7% F1 score in privacy information summarization.
2. Related Work
@@
Privacy Risks of LLMs. Recent work constructs the corpora with online personal information or user interaction history (Ishihara, 2023; Sun et al., 2023) to train LLMs (Shao et al., 2023; Song et al., 2019), and enhance inference ability (Huang et al., 2023a; Sadeq et al., 2024), resulting in the issue of memorization (Carlini et al., 2023; Smith et al., 2023; Kim et al., 2023; Carlini et al., 2021), which recalls the personal sequences from LLM corpora. Existing works solve this problem on the LLM service provider side, including data pre-processing approaches (e.g., sanitizing training data (Dernoncourt et al., 2017) and reducing the frequency of private data appearing (Kandpal et al., 2022; Lee et al., 2022)), and privacy-preserving training methods (e.g., differentially private training (He et al., 2023; Li et al., 2022; Yu et al., 2022)). In contrast, we take a different angle for privacy risks of LLMs by locally detecting privacy in the user interactions with LLMs to prevent private information from being uploaded to LLM service providers, solving the problem at the root of the user side.
@@
Privacy Protection. Most previous work focuses on privacy detection for specific documents or context shared by users on the Internet (Akiti et al., 2020; Hathurusinghe et al., 2021), thereby removing PII to achieve anonymization (Lison et al., 2021b; Papadopoulou et al., 2022; Yukhymenko et al., 2024a). However, privacy detection in user interactions with LLMs, where users have already authenticated their identifiers, is fundamentally different from these scenarios. The target application scenario of this work is to detect any newly leaked private data related to the user after a query is made, extending beyond PII. Additionally, existing approaches are typically constrained to specific topics or categories (Lison et al., 2021a; Xu et al., 2020), such as medical privacy (Zhao et al., 2019; Klein et al., 2017; Valizadeh et al., 2023, 2021), personal employment information (Tonneau et al., 2022; Preotiuc-Pietro et al., 2015), online comments (Dou et al., 2024), and code privacy (Tang and Østvold, 2024). In contrast, the types of privacy information in LLM interactions are not limited. Furthermore, most existing privacy detection models are based on discriminative models, treating the detection process as a sequence tagging task (Dou et al., 2024; Subramani et al., 2023; Herwanto et al., 2021). They can not generate specific privacy information to remind the user. MemoAnalyzer (Zhang et al., 2024) directly prompts LLMs to identify open privacy information without predefining categories, but the detection performance remains suboptimal. Besides privacy detection, PAPILLON (Li et al., 2025) introduces a method that prompts a local model to rewrite user queries before they are uploaded to prevent privacy leakage. However, its protection remains limited to PII.
Table 1. The comparison with the existing datasets.
Dataset Target Labeling Data Size Information Language
I2B2/UTHealth (Stubbs and Uzuner, 2015) Personal Health Manual 1,304 $\times$ EN
CodE ALLTAG (Eder et al., 2020) Email Manual 2,309 $\times$ DE
TAB (Pilán et al., 2022) Court Judgments Manual 1,268 $\times$ EN
Reddit-Self-Disclosures (Dou et al., 2024) Comments Manual 2,415 $\times$ EN
SynthPAI (Yukhymenko et al., 2024b) Comments Manual 7,823 $\times$ EN
Ours LLM Interactions Automated 249,683 $\surd$ EN, ZH
@@
Privacy Datasets. Existing datasets for privacy detection are limited to specific domains and there is no large-scale dataset for real-name LLM interactions. For instance, PrivacyBot (Tesfay et al., 2019) provides a sentence-level privacy classification dataset, while I2B2/UTHealth (Stubbs and Uzuner, 2015) and CEGS N-GRID (Filannino et al., 2018) focus on privacy entity recognition in the personal health domain. CodE ALLTAG (Eder et al., 2020) targets privacy entity recognition in German emails, TAB (Pilán et al., 2022) annotates private entities in court judgments, and JobStack (Jensen et al., 2021) specializes in identifying such entities in job postings. SynthPAI (Yukhymenko et al., 2024b) annotates simulated comments with inferable personal attributes, such as age and gender. Additionally, these datasets face challenges of data volume, and they are manually labeled. In contrast, our dataset focuses on LLM interactions and is annotated automatically, resulting in a significantly larger scale compared to existing datasets and encompassing a broader range of privacy. Furthermore, it offers detailed information for each extracted privacy phrase. Table 1 presents a comparison between our dataset and existing datasets.
3. Problem Formulation
@@
Privacy in LLM Interactions. We focus on the scenario where the user interacts with LLMs by submitting real-name text queries. The LLM service providers are honest-but-curious, which implies that they may attempt to infer privacy information from queries and do not actively launch attacks. Different from PII, which refers to data that can identify an individual in an anonymous text, we define privacy information in LLM interactions as the service provider’s knowledge about user-related private data or personal details, before and after the user poses a query with a real, unique identifier. Specifically, the private data in LLM interactions refers to the data that, if exposed, could negatively impact the personal, social, or professional aspects of the current user or the user’s related people, such as family, friends, or close associates. This definition of privacy information leakage is consistent with the principle of differential privacy, which considers the difference in the query output with and without each individual data record. Typical privacy information includes, but is not limited to, sensitive preferences, intentions, and opinions. We also define privacy phrases as portions of user queries that leak privacy information. The privacy phrase for the user’s first query in Figure 1 is “traveling”, and the corresponding privacy information is “The user will go traveling in the next couple of days.” We emphasize that the privacy information in LLM interactions should be related to the current user who poses the query with direct evidence. For example, as shown in Figure 1, the statement “Shanghai is also where David works.” leaks privacy information about David, but it does not leak the privacy information about the current user, since there is no direct evidence indicating David is related to the user.
@@
Figure 2. Automated privacy annotation pipeline with strong LLM and privacy detection module with light-weight LLM.
@@
Automated Privacy Annotation Requirement. As shown in Figure 2, to implement privacy leakage alerting on the user side, it is necessary to first build an automated privacy annotation pipeline based on a cloud-side LLM API to annotate dialogue datasets with privacy phrases and specific privacy information. Such a dataset can serve as an evaluation benchmark for downstream applications and provide training data for optimization. Local privacy detection methods with light-weight LLM are then designed to detect privacy from user real-time queries before they are sent to service providers, and they should be deployed locally on the user’s device or be a request proxy built by the user’s organization. Formally, we let $D$ denote the user dialogue with LLM and $q$ denote a query from $D$. The goal is to extract the privacy phrase $\mathcal{P}_{q}=\{p_{1},...,p_{m}\}$ and summarize privacy information $\mathcal{I}_{q}=\{I_{1},...,I_{m}\}$, where $m$ is the number of the extracted phrases, $p_{j}$ is a segment of $q$ that discloses private or sensitive information about the user, and $I_{j}$ is a statement summarizing the specific details revealed by $p_{j}$.
@@
Figure 3. Overview of our automated pipeline to extract privacy phrase and annotate privacy information over dialogue datasets: ➊ determining whether the input leaks privacy; ➋ extracting privacy categories from user queries that leaks privacy; ➌ extracting privacy phrases from user queries with the help of privacy categories; and ➍ annotating privacy information for each phrase.
4. Automated Privacy Annotation Pipeline for Dialogue Datasets
We design an automated four-step pipeline based on the strong LLM API to extract privacy phrases and annotate privacy information for dialogue datasets.
4.1. Pipeline Rationale
A straightforward approach to extract privacy information is directly prompting GPT-4o. However, this method may force GPT-4o to extract information from samples that do not contain any privacy leaks, which results in a decrease in precision. Furthermore, the original understanding of privacy by GPT-4o revolves around the idea of whether a privacy phrase can reveal a user’s identifiable information. While this is intuitive for identifying PII, it does not align with our intended focus in Section 3, which aims to detect privacy information in real-name LLM interactions. We also observe that the results of direct prompting are limited to a few common types of privacy, such as gender, with less accuracy in detecting more nuanced aspects like user status or intent.
To address these limitations, we developed the privacy annotation pipeline, as shown in Figure 3. (1) To avoid privacy extraction for user dialogue that does not involve privacy leakage, the first step is a filter, classifying dialogues into two categories: leaked and non-leaked. (2) Since the diverse and open-ended nature of privacy categories in user interactions with LLMs, the second step aims to automatically collect a comprehensive coverage of privacy categories for the corpus. We independently extract privacy categories for each sample and take the union of all categories as the extensive set of privacy categories at the dataset level. (3) Then, with the assistance of the extensive category set, we extract privacy phrases for each sample based on the whole context of queries and filter the results based on the rules of being user-related and having a clear reference. (4) The last step is to annotate privacy information for each extracted phrase. We also compare the performance of direct prompting with our automated pipeline in Section 4.3. All associated prompts can be found in Appendix A.
4.2. Pipeline Design
Privacy Leakage or Not Classification. We explicitly address the scenario where the user interacts with identifier exposure, and prompt GPT-4o to assess whether the user has disclosed any new privacy information during the interaction. We further clarify that privacy information must be linked to the user or individuals associated with them, and it encompasses various aspects such as opinions, preferences, and personal details. To ensure more reliable inferences, we not only ask GPT-4o to generate a simple yes or no judgment, but also request it to explain the reasoning before its decision. Additionally, we provide a set of manually annotated samples as a reference to guide the model.